whats the sum of the measures of the exterior angle of an octogan?
1
Expert's answer
2012-10-25T11:50:32-0400
interior sum =pi ( n-2)
n=8
pi*6
exterior = pi - inetrior
sum
exterior=n*pi-sum interior
sum exterior= 8*pi-6*pi=2pi